In 3 years,alex will be 3 times as old as his sister Precy.A year ago,Alex was 7 times as old as Precy.How old are they now?
Answer by alka001(14)

Let Precys present age be x yrs
A year ago her age was (x-1) yrs
So Alex age was 7(x-1) yrs
Alex present age is 7(x-1)+1
In 3 yrs Precys age will be x+3
In 3 yrs Alex age will be 3(x+3)
3(x+3)=7(x-1)+1+3
3x+9=7x-7+4
3x-7x=-11-9
-4x=-20
x=5
So Precys present age is 5 yrs and Alex present age is 7(5-1)+1=28+1=29 yrs